Find curvature r(t) = ( ½ Cost, 1-sint, -3/2 cost)

1
Expert's answer
2022-01-18T06:15:55-0500
"r(t)=\langle\dfrac{1}{2}\cos t, 1-\sin t, -\dfrac{3}{2}\cos t\rangle"

"r'(t)=\langle-\dfrac{1}{2}\sin t, -\cos t, \dfrac{3}{2}\sin t\rangle"

"r''(t)=\langle-\dfrac{1}{2}\cos t, \sin t, \dfrac{3}{2}\cos t\rangle"

"|r'(t)|=\sqrt{(-\dfrac{1}{2}\sin t)^2+( -\cos t)^2+(\dfrac{3}{2}\sin t)^2}"

"=\dfrac{\sqrt{10-6\cos ^2t}}{2}"

"r'(t)\times r''(t)=\begin{vmatrix}
 i & j & k \\
\\
 -\dfrac{1}{2}\sin t & -\cos t &\dfrac{3}{2}\sin t \\ \\
 -\dfrac{1}{2}\cos t & \sin t &\dfrac{3}{2}\cos t
\end{vmatrix}"

"=i\begin{vmatrix}
 -\cos t & \dfrac{3}{2}\sin t \\ \\
 \sin t & \dfrac{3}{2}\cos t
\end{vmatrix}-j\begin{vmatrix}
 -\dfrac{1}{2}\sin t & \dfrac{3}{2}\sin t \\ \\
 -\dfrac{1}{2}\cos t& \dfrac{3}{2}\cos t
\end{vmatrix}"

"+k\begin{vmatrix}
 -\dfrac{1}{2}\sin t & -\cos t \\ \\
 -\dfrac{1}{2}\cos t & \sin t
\end{vmatrix}=-\dfrac{3}{2}i-\dfrac{1}{2}k"

"|r'(t)\times r''(t)|=\sqrt{(-\dfrac{3}{2})^2+(0)^2+(-\dfrac{1}{2})^2}=\dfrac{\sqrt{10}}{2}"

Find curvature


"\kappa(t)=\dfrac{|r'(t)\times r''(t)|}{(|r'(t)|)^{3}}"

"=\dfrac{\dfrac{\sqrt{10}}{2}}{(\dfrac{\sqrt{10-6\cos ^2t}}{2})^{3}}"

"=\dfrac{4\sqrt{10}}{(10-6\cos ^2t)^{3/2}}"

"=\dfrac{2\sqrt{5}}{(5-3\cos ^2t)^{3/2}}"

"\kappa(t)=\dfrac{2\sqrt{5}}{(5-3\cos ^2t)^{3/2}}"